An evacuated reaction vessel is filled with NOBr to an initial pressure of 5.00 atm. When the system reaches equilibrium according to the reaction below, there are 0.750 atm of Br₂. What is Kp for this reaction?


2 NOBr (g) ⇌ 2 NO (g) + Br₂ (g)


1
Expert's answer
2021-03-22T06:32:01-0400

Consider this reaction process as shown below:

Reaction : "2 NOBr (g) \u21cc 2 NO (g) + Br\u2082 (g)"

At "\ t=0\ \ \ \ \ \ \" "5\ \ \ \ \ \ \ \" "\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0\ \ \ \ \ \ \ \ \ \ \ \ \ \"

At "\ t=T""\ \ \ \ \ 5-2\alpha\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 2\alpha \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \alpha"

And, at equilibrium time ,pressure of "Br_2" is"\ \alpha =0.750\ atm"

So ",\ p_{NO}=2\alpha=2\times 0.750=1.5\ atm"

And,"p_{NOBr}=5-2\times 0.750=5-1.5=3.5\ atm"

"K_p=\frac{[p_{NO}]^2[p_{Br_2}]}{[p_{NOBr}]^2}" "=\frac{1.5^2\times 0.750}{3.5^2}\ atm=0.1377\ atm\approx0.14\ atm"